2. The Capital structure of ABC Ltd, is as under:
Equity share capital - ₹ 100 Lacs
10% Debentures - ₹ 50 Lacs

• The sales for the year 2019 are 1.5 Lac units@ ₹ 40per unit
• Also, the variable cost per unit is 20 % of sales revenue
• ₹ 12 Lacs is the fixed operating cost.
• Assume Income tax rate as 40 %

Calculate Operating, Financial and Combined Leverage of the firm and interpret the result.
1
Expert's answer
2020-06-11T10:56:21-0400

Operating Leverage

"OL=\frac{S-VC}{FC}=\frac{60-12}{12}=4"


Financial Leverage

"FL=\frac{ Debentures }{Equity share capital }=\frac{50}{100}=0.5"


Combined Leverage

"CL=\frac{60-12}{60-12-12}=1.3"


greatest effect print Operating Leverage